OpenStudy (ray1998):

Please help! Will give fan and medal. When treated with an antibiotic 94% of all dolphins are cured of an ear infection. If 6 dolphins are treated, find the probability that exactly 4 are cured. The probability is _______. Thank you in advance :)

OpenStudy (anonymous):

4/6 = 2/3 = .66 or 67%

OpenStudy (ray1998):

That's all it wants you to do?

OpenStudy (ray1998):

Okay.. well thank you very much for your time and effort!

OpenStudy (anonymous):

yes, no problem